Technical Observing - 20131025 UT

Observers: JHill, DMiller (Tucson)
Telescope Operator: SAllanson (LBTO)
AO Support:: DMiller (Tucson)
Telescope Support:: MMidkiff (LBTO)
Instrument Support:: JPower (LBTO)
Engineering Support: DAshby (Tucson)


Lost 2 hours at beginning for Clouds and VFD problems in Utility Chiller #2.

Had strange problems with telescope tracking drifting fast on the left side, but not the right. Also had time problems on jet (Issue 4880). We suspected the hexapod attached to rigid secondary, but it checked OK. Drift went away when we restarted PCS. ??

Standard recipe for adjusting the remounted SX M2 did not work, and ran us into limits. Resorted to measuring binodal aberration from scratch and adjusting M2 position. We successfully aligned Rigid M2, but have not yet understood why the standard recipe didn't work. MODS1 and LUCI1 are ready for science with the Rigid Secondary.

Found a bug in the start of guiding when a single asyncronous preset is sent in binocular mode. See Issue 4881.

We adjusted the AGw2 probe onto the axis of rotation within ~0.4 arcsec.

We verified the anomalous pupil and guide star wobbles in AGw2 has been cured by replacing the mirror. The upper limit to the present wobble is 0.2 arcsec (10x improved from a month ago).

We measured the AGw2 pupil wobble (data needs more analysis in daytime).

We adjusted the binodal astigmatism at RFBG. This completes the AGw2 alignment until LUCI2 is available for focus and transformation measurements.

We used the special version of PSF to measure Z4-Z11-Z22 cross-talk at LFBG and RFBG.

We took Z4-Z11-Z22 cross-talk data on both sides with the special version of PSF.


Closed for threatening clouds at the beginning. Some VFD problems on Utility Chiller 2 at the same time.


02:45 Open

Initial Re-Alignment for Rigid Secondary with LUCI@LFBG

02:54 ACQUIRE preset in PARALLACTIC mode

Fri Oct 25 02:54:29.061 2013 Preset alert for left LUCIFER mode ACQUIRE: Invalid buffer: RPCERROR

AOSL died, resend preset.

02:58 Found no star on left at the default IE=-53 CA=-52 (as expected). Spiral to find the star and focus it. EL=69

03:03 Adjust Tip-Tilt to get back to original IE/CA Tip=-19 Tilt=+62

03:08 Adjust Coma on M2 by sending Z7 and Z8.

We are near the RX Limit of M2. Are the assumptions of this procedure (in BGTechProcedures) still valid when we have changed hexapods?

Active X=-0.745 Y=-6.056 Z= RX=+631 RY=-77.6
Global X=-3.21 Y=+0.102 Z=-3.8 RX=-42 RY=-569

LUCI on both sides.

AOS can't find LUCI flat file on right side. Doug will make the link from the latest flat.

03:28 ACQUIRE preset on both sides, BUT SX M2 runs into a radius limit (dominated by X=-8.4 mm).

Back to LUCI@RFBG in monocular mode

Steve notes that maximum latency on MCSPU is up to 675 ms. Could our large CA correction have a contributing time error? Called TomS

03:39 Recenter star with Tip-Tilt, but this drives us into the RY limit.

LUCI on both sides

Now we abandon the traditional realignment recipe.
Remove Tip-Tilt from SX M2.
Clear the Coma Correction on SX M2.

03:48 ACQUIRE Preset on both sides - See star on right, but not on left as expected from above

03:53 Change SX Xglobal from -3.2 to 0 mm. Moving total X up to -4.6 mm.

04:00 Adjusting Coma with M2 XY on Left

SX M2 Global X=-3.00 Y=-2.00 Z=-3.8 RX=-42 RY=-569
SX IE=-57 CA=-80
DX IE=-4 CA=-199

DanC has sync'd up the DSPs and MCSPU (with no change in pointing). Apparently jet had disconnected from NTP - see Issue 4880.

Attempting Field Data BS9188

04:09 ACTIVE on Left, ACQUIRE on Right at EL=82

04:12 ACTIVE async on Right - Guiding on Right is oscillating by +-4 arcsec. Why?

Restart GCSR

04:15 Resend ACTIVE async on Right - still oscillates

04:17 stop reporting centroids on Right until Active Optics stabilizes

04:23 Resume guiding while Doug adjusts hotspot to 266,249. Saved in

04:27 Guiding on Left is horribly systematic, while guiding on right now seems fine.

Called Dave A. He doesn't have any other suspects beyond hexapod firmware. The plan is to restart a bunch of things, and then measure the angular scaling of both sides with tip-tilt.

Restart GCSL

04:46 ACTIVE on Left - wfs image doesn't arrive, but recovered on its own.

Doug has checked pupil centering and hotspot on left.

Now apparent drift is in different direction.

04:51 Restart PSFL and PSFR while the preset is running.

04:53 resend ACTIVE presets to both sides, but missing new global offsets on left.

04:55 resend ACTIVE async on left - but coma and pointing issues

Seeing on guiders is 0.35 and 0.40 arcsec FWHM respectively (L & R).

05:02 Restart PCS

05:05 Send ACQUIRE sync presets in PARALLACTIC mode to both sides (still BS9188)

Right IE=+9 CA=-70

Spiral both to find star on Left at IE=-116 CA=-70

05:14 Send Tip = +10 to each M2 sequentially. Motions are indistinguishably the same on the two sides.

05:16 Send Tilt = +10 to each M2 sequentially - motions are apparently identical within the 5% error of your eye. Directions were cleanly orthogonal to the parallactic axes.

05:19 ACTIVE sync presets on-axis on both sides.

Now guiding is great on both sides -- 0.07 arcsec rms -- as good as it ever gets.
We DO NOT understand what happened.
Was the NTP issue messing up the pointing in azimuth? 0.7 sec of time (10 sec of angular mispointing) doesn't seem to be enough to explain our problems with either drift or M2 collimation.
Was PCS confused in some way? The circumstantial evidence is that restarting PCS cured the problem.
We can't see any problem with either hexapod as measured by applying 10 arcsec Tip-Tilt (coma-free pointing) corrections on-sky in ACQUIRE mode.

Field Data on LFBG for BS9188 in Binocular Mode

05:33 Preset on Left for the beginning of Field Aberration Data. This left async preset made a 0.3 arcsec jump on right that guiding corrected.

05:40 Preset on Left on-axis at different PA - left side is oscillating by +-2 arcsec. Stopped it by not reporting centroids for ~10 sec.

Doug is fighting with a problem in the IDL code.

05:51 Preset sync on left, oscillates +-3 arcsec. See Issue 4881.

LUCI1@LFBG Field Data on BS9188

data in directory /home/lbto/data/20131025

05:55 Switch to monocular mode on left

05:57 The acquisition offset for this preset was elongated (left acquire 000011) and the first guide image was trailed (left guide 000750). Is the oscillation problem above caused by something other than binocular mode? Or was this an after effect of binocular mode?

06:01 Preset ACTIVE on left. Acquisition was rock steady.

06:10 Rotate DX M3 to LN position with OSSGUI.

06:20 Cube of 100 x 1 sec images for sky background without collimating anything.

06:22 Doug wants Tip=-40, and dIE=+40 (~2/3 of his measurement). Tilt=-70 and dCA=+70
SX Total Collimation after that adjustment is X=-0.941 Y=-2.417 Z=-2.225 RX=+231 RY=+365

06:28 Cube of 100 x 1 sec after collimating M1/M2 on right.

06:31 Tip 5 0 on Right (M1M2) Cube of 100 x 1 sec

These irtc1 cubes are in /newdata with headers only from the IRTC GUI (not taken through LBTtools.observe.takepic). There were taken with the empty filter, and the focus position was not compensated.

06:55 Doug wants Tip=-80 and dIE=+40, and Tilt=-30 and dCA=-40. IE=-36 CA=-40.

07:24 SX IE=-49 CA=-40 saved in SX_Gregorian.20131025.ptmod

07:25 Final Rigid Secondary Collimation Values
SX Total@EL=54.7 X=-3.600 Y=+0.240 Z=-2.517 RX=-160 RY=-18
SX Global X=-3.000 Y=-2.000 Z=-3.800 RX=-42 RY=-569.2
SX Manual P X=+1.532 Y=4.085 Z=0.000 RX=-785.6 RY=+294.6
SX Active X=-1.458 Y=-2.958 Z=0 RX=+308 RY=-149
SX NewGlobal X=-2.926 Y=-0.872 Z=-3.800 RX=-519.6 RY=-423.6

Installed these new Global Offsets in SXSMLUCIFERCollimation.dat. Previous version was SXSMLUCIFERCollimation.20130521.dat.
Installed these new Global Offsets in SXSMMODSCollimation.dat. Previous version was SXSMMODSCollimation.20130621.dat.
Saved copy in SXSMLUCIFERCollimation.20131025rigid.dat

AGw2 Alignment at RFBG

data in directory /home/lbto/data/20131025/agw2

07:32 Authorize LUCI2@RFBG

07:34 ACTIVE preset on Right

07:42 0002 - was at 0,619.5 not 0,612.5

setxy -u 2 -x 0 -y 612.5

07:45 0003 - nice but too large

07:48 0004 - X=166.5 Y=104.5 ===> 333+17, 209+15 ===> 350,224 where we want to match hotspot of 266,249
need Y decreased by 25 pixels = 1.25 arcsec = 750 microns so increase RAD ooffset by 750 microns ===> 419147 to 419897
need X decreased by 84 pixels = 4.2 arsec = 24xx microns so increase ROT homeoffset by 68 x 2.4 = 163 ===> -33876 to -33713
departure from circularity is 0.2 arcsec (improved ~10x from before the mirror replacement)

08:10 stop and restart GCS, AGW2, oacserver

08:16 send ACTIVE preset on Right to BS9188 seeing is 0.5 arcsec at EL=45

Stop guiding, take fake images, move probe to 0,612.5, start rotator, readGuideCam

08:19 0005 - X=127.5 Y=92.5 ===> 255+17, 185+15 ===> 272,200 decrease ooffset by 375, ooffset = 419522, and increase by 12, homeoffset = -33701
Made rotation limits more negative by 175 ===> lim_f=-13339 lim_n=13041

08:32 stop and restart GCS, AGW2, oacserver

08:37 ACTIVE preset BS9107 on-axis at EL=85

Why did the guide probe still go to 0, 619.75 ? I don't understand this at high elevation where refraction is small.

08:46 0006 - X=127.5 Y=96.5 ===> 255+17,193+15 ===> 272,208
So we are off by 41 pixels in Y which is 2 arcsec.

Pupil Wobble measurement for AGw2

08:56 ACTIVE preset on right

Measured a wobble of 3 pixels (even with the wrong RIGHTPOINTINGORIGIN).

09:22 edit PCSInstrument.conf to remove RIGHTPOINTINGORIGIN of 0,-7.7 mm. That was why probe was going to 619.7. This was an artifact of the PISCES campaign.

09:23 Stop and restart PCS to get the new PCSInstrument.conf

Failed to find star - the one it grabbed was too faint (because we forgot to restore IE/CA, and RIGHTPOINTINGORIGIN changed).

09:28 ACQUIRE on WT10_342 IE=8 CA=-193

09:29 ACTIVE preset back to BS9107

stop WFSing

data in directory /home/lbto/data/20131025/pupil_wobble_right_2/


Field Data for AGw2@RFBG


10:26 Adjust for DX M2 Tip=+40 dIE=-40 and Tilt +10 dCA=-10


10:30 Preset cancelled for rotator off-source


11:09 Adjust for DX M2 Tip=+30 dIE=+10 and Tilt +10 dCA=0

11:12 field_collect at EL=58

DX IE=-22 CA=-203 saved in DX_Gregorian.20131025.ptmod

DX M2 Globals are zeros
DX M2 Manual pointing X=-0.511 Y=-1.532 Z=0.000 RX=+294.6 RY=-98.2
DX M2 Active X=+0.603 Y=+2.707 Z=0.071 RX=-282.1 RY=+59.6
DX NewGlobal X=+0.092 Y=+1.175 Z=+0.071 RX=+12.5 RY=-38.6

These Global Offsets were applied to DX M2 for all Gregorian focal stations. DXSMLUCIFERCollimation.20131025.dat

Changed link for PSF executable from PSF.20130919 to PSF.20131024b.

M1 Zernike Cross-talk data on Both sides

11:59 Restart PCS for new PCSInstrument.conf

12:00 Authorize LUCI on both sides

12:xx ACTIVE presets

12:xx GUIDE presets

12:05 zernike_collect

12:30 Why is left side 4x more noisy on Z5-Z11? Doug is having trouble keeping SX collimated.

13:05 Restore the PSF link to the normal PSF.20130919 version. Steve will restart them momentarily.


13:07 Close for impending dawn with 0.7 arcsec seeing at EL=37 deg!

-- JohnHill - 25 Oct 2013
Topic revision: r2 - 25 Oct 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback